OpenStudy (unklerhaukus):

Laplace transform of \[g(t) =t^2\sinh(3t)\] \[G(p) =\mathcal L\big\{t^2\sinh(3t)\big\}\]

OpenStudy (unklerhaukus):

is it best to use \[\mathcal L\{tf(t)\}=-\frac{d}{dx}\mathcal L\{f(t)\}\] a couple of times?

OpenStudy (unklerhaukus):

and then the quotient rule?

OpenStudy (unklerhaukus):

Or instead to substitute in the exponential identity for the hyperbolic trig function

OpenStudy (anonymous):

Does this approach (see attached) make sense to you?

OpenStudy (anonymous):

i think the solution maybe (1/((p-3)^3)-(1/((p+3)^3)

OpenStudy (unklerhaukus):

yeah that method works

OpenStudy (unklerhaukus):

is there some way to simplify the numerator?

OpenStudy (anonymous):

I have not tried, too tedious

OpenStudy (unklerhaukus):

i had this working

OpenStudy (unklerhaukus):

\[18(s^2+3)=(s+3)^3-(s-3^3)\] ?

OpenStudy (anonymous):

I think best expression is:\[\frac{ 1 }{ (s-3)^3 }-\frac{ 1 }{ (s+3)^3 }\]

OpenStudy (anonymous):

@unkleRhaukus:ya...i also followed the same method.

OpenStudy (unklerhaukus):

how can i get my answer to equal that nice expression ?

OpenStudy (anonymous):

if u use partial fractions..u can get the expression got by carlosGP!

OpenStudy (unklerhaukus):

ill try that now @susanka

OpenStudy (unklerhaukus):

i got \[\frac{p^2+3}{(p^2-9)^3}=\frac1{(p^2-9)^2}+\frac3{(p^2-9)^3}\]

OpenStudy (anonymous):

gr8:)

OpenStudy (chihiroasleaf):

I also get the same as @CarlosGP , I substitute into exponential identity for the hyperbolic function

OpenStudy (anonymous):

@UnkleRhaukus You must solve: \[\frac{ 18p^2+54 }{ (p^2-9)^3 }=\frac{ A }{ (p-3) }+\frac{ B }{ (p-3)^2 }+\frac{C }{ (p-3)^3 }+\frac{ D }{ (p+3) }+\frac{ E }{ (p+3)^2 }+\frac{ F }{ (p+3)^3 }\]To get C=1 and F=-1 (A=B=D=E=0)

OpenStudy (unklerhaukus):

oh, i though it was \[\frac{p^2+3}{(p^2-9)^3}=\frac{Ap+B}{p^2-9}+\frac{Cp+D}{(p^2-9)^2}+\frac{Ep+F}{(p^2-9)^3}\]

OpenStudy (anonymous):

ya..but carlosGP is right..

OpenStudy (anonymous):

@UnkleRhaukus with your formula, what happens with the p^2-9 of the denominator? it can also be factorized

OpenStudy (unklerhaukus):

im not very good with cubics

OpenStudy (chihiroasleaf):

it seems it's easier by substitute the hyperbolic function by exponential function ... :D

OpenStudy (anonymous):

Never mind! You made me hesitate. That means I was not very convinced either

OpenStudy (unklerhaukus):

where'd you get that partial fractions substitution @CarlosGP it doesn't match with my table

OpenStudy (anonymous):

Second case

OpenStudy (unklerhaukus):

but there is a p^2 in the denominator

OpenStudy (anonymous):

True, but still can be factorized:\[(p^2-9)^3=(p-3)^3(p+3)^3\] You have applied fourth case that like third case is applied when polynomial ax^2+bx+c has no real solutions and cannot be factorized. In our case we have a polynomial with roots 3 and -3 with multiplicity =3